# Normal Scalar Curvature Inequality on a Class of Austere Submanifolds

### v1

@@ Abstract.
J. Q. Ge is partially supported by NSFC (No. 12171037) and the Fundamental Research Funds for the Central Universities.
Y. Zhou is partially supported by NSFC (No. 12171037, 12271040), China Postdoctoral Science Foundation (No. BX20230018) and National Key R&\&&D Program of China 2020YFA0712800.
@@ 1. Introduction
for $n\times n$ real symmetric matrices $B_{1},\ldots,B_{m}$, where $[A,B]:=AB-BA$ denotes the commutator and $\|B\|^{2}:=\operatorname{tr}(BB^{t})$ denotes the squared Frobenius norm. After some partial results [9, 14, 15, 33], the DDVV inequality (1.2) was finally proved by Lu [34] and Ge-Tang [21] independently. In addition, the equality condition in (1.2) were completely determined in [21]. Hence we know that the equality in (1.1) holds at some point $p\in M$ if and only if there exist an orthonormal basis $\{e_{1},\cdots,e_{n}\}$ of $T_{p}M$ and an orthonormal basis $\{\xi_{1},\cdots,\xi_{m}\}$ of $T_{p}^{\bot}M$ such that the matrices corresponding to the second fundamental form along directions $\xi_{1},\cdots,\xi_{m}$ are
@@
where $P\in\operatorname{O}(n)$ and $R=(R_{jk})\in\operatorname{O}(m)$ acts as a rotation on the matrix tuple $(P^{t}B_{1}P,\cdots,P^{t}B_{m}P)$. In addition, we use the notation $E_{ij}$ to denote the $n\times n$ matrix with $1$ in position $(i,j)$ and 00 elsewhere.
@@
In order to illustrate that the above theorems make sense, we give some explicit examples of austere submanifold of type $\mathcal{Q}(p,q)$, including Bryant’s generalized helicoids, two austere embeddings of Grassmann manifolds $\operatorname{G}_{2}(\mathbb{R}^{4})$ and $\operatorname{G}_{2}(\mathbb{C}^{4})$, and some isoparametric focal submanifolds satisfying Condition A in unit spheres. In fact, among focal submanifolds of isoparametric hypersurfaces with $4$ distinct principal curvatures, we find that the focal submanifolds $M_{+}$ with multiplicities $(1,2k)$ and the focal submanifold $M_{-}$ with multiplicities $(4,3)$(definite) are austere submanifolds of type $\mathcal{Q}(p,q)$ for some suitable integers $p,q$. It is worth pointing out that the focal submanifold $M_{+}$ with multiplicities $(1,2)$ achieves the equality in Theorem 1.3.
@@ 2.1. Preparatory lemmas
where $E_{ij}$ is the $n\times n$ matrix with $1$ in position $(i,j)$ and 00 elsewhere. Then $\{\hat{E}_{ij}\}$ is an orthonormal basis of $\mathcal{Q}(p,q)$. Next, we put an order on the index set $S$ by
(2.1) $(i,j)<(k,l)\ \text{if and only if}\ i<k\ \text{or}\ i=k,j<l,$
and use this order to identify an index $(i,j)\in S$ with a single (Greek) index $\alpha\in\{1,...,N\}$, i.e., $(i,j)\ \hat{=}\ \alpha=(i-1)q+(j-p)$.
@@
$[P_{2}^{t}\hat{Q}_{\alpha}P_{2},P_{2}^{t}\hat{Q}_{\beta}P_{2}]=\begin{pmatrix}\left[D,X_{\beta}\right]&\frac{1}{\sqrt{2}}Dc^{t}_{\beta}&DY^{t}_{\beta}+Y^{t}_{\beta}D\\ -\frac{1}{\sqrt{2}}c_{\beta}D&0&-\frac{1}{\sqrt{2}}c_{\beta}D\\ DY^{t}_{\beta}+Y^{t}_{\beta}D&\frac{1}{\sqrt{2}}Dc^{t}_{\beta}&\left[D,X_{\beta}\right]\\ \end{pmatrix}$
@@
Let $C(E)$ denote the matrix in $M(\binom{N}{2})$ defined by
@@
(2.15) $C(B)=\varphi(B^{t})C(E)\varphi(B),\hskip 8.5359ptC(Q)=\varphi(Q^{t})C(E)\varphi(Q).$
@@
where +$\mathbb{R}^{N}_{+}:=\left\{x=\left(x_{1},\cdots,x_{N}\right)\in\mathbb{R}^{N}\setminus\{0\}:x_{\alpha}\geq 0\ \mbox{for any}\ 1\leq\alpha\leq N\right\}$.
@@
With the given prerequisites, then $y$ is a maximum point of $f_{Q}\left(x\right)$ in the cone spanned by $\Delta_{\varepsilon}$ and an interior maximum point of $f_{Q}\left(x\right)$ in $\Delta_{\varepsilon}^{\gamma}$. Hence, applying the Lagrange Multiplier Method, there exist numbers $b_{\gamma+1},\cdots,b_{N}$ and a number $a$ such that
@@
(2.21) $\sum_{\beta=1}^{N}y_{\beta}\|[\hat{Q}_{\alpha},\hat{Q}_{\beta}]\|^{2}-\frac{1}{2}=\Big{\{}\begin{array}[]{ll}a&\alpha\leq\gamma,\\ b_{\alpha}&\alpha>\gamma.\end{array}$
@@
Hence $S\setminus(J\cup\{1\})\neq\varnothing$, and the last $\leq$” in (2.24) should be “$<$” by the definition of $J$ and the positivity of $y_{\beta}$ for $\beta\in S\setminus(J\cup\{1\})$. $\Box$
@@
$y\in\Delta^{\gamma}:=\left\{x\in\Delta:x_{\alpha}>0\ \mbox{for}\ \alpha\leq\gamma,\ \mbox{and}\ x_{\beta}=0\ \mbox{for}\ \beta>\gamma\right\},$
@@
where $D=\operatorname{diag}(\lambda_{1},\ldots,\lambda_{q})$. In the proof of Lemma 2.3 with $J_{\alpha}=J$, the first line in (2.9) shows that $c_{\beta}=0$, $b_{\beta}$ is skew-symmetric and $b^{\beta}_{ij}=-b^{\beta}_{ji}=0$ for all $\beta\in J$ and $(i,j)\in\{(i,j):1\leq i<j\leq q\}\setminus I$; the third line in (2.9) shows that $b^{\beta}_{ij}=0$ for all $\beta\in S\setminus(J\cup\{1\})$ and $(i,j)\in I$. For the eigenvalues ±$\pm\lambda_{1},\ldots,\pm\lambda_{q}$ of $\hat{Q}_{1}$, because of the equality condition of Lemma 2.2, we have to consider the following cases:
@@ 3.3. Some isoparametric focal submanifolds satisfying Condition A
For an isoparametric hypersurface in a unit sphere with multiplicities $(m_{1},m_{2})$, there exist two focal manifolds $M_{+}$ and $M_{-}$ of codimension $m_{1}+1$ and $m_{2}+1$, respectively. A point in a focal manifold is said to satisfy Condition A if the kernels of all shape operators coincide [36]. The Condition A plays an important role in the research of isoparametric hypersurfaces in unit spheres with $4$ distinct principal curvatures. In [23], the subset $C_{A}$ of the points satisfying Condition A was determined for all the focal submanifolds. Thus, we know that the focal submanifolds satisfying Condition A everywhere are $M_{+}$ with multiplicities $(2,2)$, $(1,k)$($k\in\mathbb{N}^{+}$) and $M_{-}$ with multiplicities $(2,1)$, $(6,1)$, $(4,3)$(definite). We are going to find austere submanifolds of type $\mathcal{Q}(p,q)$ in these focal submanifolds.
@@
$m$ $1$ $2$ $3$ $4$ $5$ $6$ $7$ $8$ $\cdots~{}m+8$
$\delta(m)$ $1$ $2$ $4$ $4$ $8$ $8$ $8$ $8$ $\cdots~{}16\delta(m)$
It follows that the focal submanifolds $M_{+}$ with multiplicities $(1,2k)$ are austere submanifolds of type $\mathcal{Q}(m_{2},m_{1}+m_{2})$. By interchanging the multiplicities $(m_{1},m_{2})$, we also obtain that the focal submanifolds $M_{-}$ with multiplicities $(2,1)$, $(6,1)$, $(4,3)$(definite) are austere submanifolds of type $\mathcal{Q}(m_{1},m_{1}+m_{2})$. However, the focal submanifolds $M_{-}$ with multiplicities $(2,1)$, $(6,1)$ are congruent to focal submanifolds $M_{+}$ with multiplicities $(1,2)$, $(1,6)$.
@@
Let $M_{+}$ ((((resp. $M_{-}$)))) be the focal submanifold of an isoparametric hypersurface with $4$ distinct principal curvatures and codimension $m_{1}+1$ ((((resp. $m_{2}+1$)))) in a unit sphere.
@@
    The focal submanifold $M_{-}$ with multiplicities $(4,3)$((((definite)))) is an austere submanifold of type $\mathcal{Q}(4,7)$.
@@ 4.2. Simons-type gap theorem
Since $\theta_{1},\cdots,\theta_{n}$ are orthonormal, we have $\theta_{j}=0$ for $j\geq 5$, and thus $\operatorname{dim}M=4$. Direct calculations show that for any $a,b\in\mathbb{R}$, the eigenvalues of $aC_{1}+bC_{2}$ are ±$\pm\sqrt{a^{2}+b^{2}}$ of multiplicity 2. In other words, $M$ has at most two distinct principal curvatures of multiplicity $2$ in every normal direction. Then it follows from [30, Theorem 1.5] that $M$ is given by the Veronese embeddings of complex projective plane. But the matrices corresponding to the second fundamental form of this Veronese embedding does not satisfy condition (4.4). (One can find an explicit description in [32, p. 560].) ∎

### v2

@@ Abstract.
J. Q. Ge is partially supported by NSFC (No. 12171037, 12571049) and the Fundamental Research Funds for the Central Universities.
Y. Zhou is partially supported by NSFC (No. 12171037, 12271040), China Postdoctoral Science Foundation (No. BX20230018) and National Key R&\&D Program of China 2020YFA0712800.
@@ 1. Introduction
for $n\times n$ real symmetric matrices $B_{1},\ldots,B_{m}$, where [A,B]:=ABBA[A,B]:=AB-BA denotes the commutator and $\|B\|^{2}:=\operatorname{tr}(BB^{t})$ denotes the squared Frobenius norm. After some partial results [9, 14, 15, 33], the DDVV inequality (1.2) was finally proved by Lu [34] and Ge-Tang [21] independently. In addition, the equality condition in (1.2) were completely determined in [21]. Hence we know that the equality in (1.1) holds at some point $p\in M$ if and only if there exist an orthonormal basis $\{e_{1},\cdots,e_{n}\}$ of $T_{p}M$ and an orthonormal basis $\{\xi_{1},\cdots,\xi_{m}\}$ of $T_{p}^{\bot}M$ such that the matrices corresponding to the second fundamental form along directions $\xi_{1},\cdots,\xi_{m}$ are
@@
where $P\in\operatorname{O}(n)$ and $R=(R_{jk})\in\operatorname{O}(m)$ acts as a rotation on the matrix tuple $(P^{t}B_{1}P,\cdots,P^{t}B_{m}P)$. In addition, we use the notation $E_{ij}$ to denote the $n\times n$ matrix with $1$ in position $(i,j)$ and 0 elsewhere.
@@
In order to illustrate that the above theorems make sense, we give some explicit examples of austere submanifold of type $\mathcal{Q}(p,q)$, including Bryant’s generalized helicoids, two austere embeddings of Grassmann manifolds $\operatorname{G}_{2}(\mathbb{R}^{4})$ and $\operatorname{G}_{2}(\mathbb{C}^{4})$, and some isoparametric focal submanifolds satisfying Condition A in unit spheres. In fact, among focal submanifolds of isoparametric hypersurfaces with $4$ distinct principal curvatures, we find that the focal submanifolds $M_{+}$ with multiplicities $(1,2k)$ and the focal submanifold $M_{-}$ with multiplicities (4,3)(4,3)(definite) are austere submanifolds of type $\mathcal{Q}(p,q)$ for some suitable integers $p,q$. It is worth pointing out that the focal submanifold $M_{+}$ with multiplicities $(1,2)$ achieves the equality in Theorem 1.3.
@@ 2.1. Preparatory lemmas
where $E_{ij}$ is the $n\times n$ matrix with $1$ in position $(i,j)$ and 0 elsewhere. Then $\{\hat{E}_{ij}\}$ is an orthonormal basis of $\mathcal{Q}(p,q)$. Next, we put an order on the index set $S$ by
(2.1) (i,j)<(k,l)if and only ifi<kori=k,j<l,(i,j)<(k,l)\ $\text${if and only if}\ i<k\ $\text${or}\ i=k,j<l,
[…]
@@
$P_{2}^{t}\hat{Q}_{\alpha}P_{2},P_{2}^{t}\hat{Q}_{\beta}P_{2}]=\begin{pmatrix}\left[D,X_{\beta}\right]&\frac{1}{\sqrt{2}}Dc^{t}_{\beta}&DY^{t}_{\beta}+Y^{t}_{\beta}D\\ -\frac{1}{\sqrt{2}}c_{\beta}D&0&-\frac{1}{\sqrt{2}}c_{\beta}D\\ DY^{t}_{\beta}+Y^{t}_{\beta}D&\frac{1}{\sqrt{2}}Dc^{t}_{\beta}&\left[D,X_{\beta}\right]\\ \end{pmatrix}$
@@
Let $C(E)$ denote the $\binom{N}{2}\times\binom{N}{2}$ matrix defined by
@@
(2.15) $C(B)=\varphi(B^{t})C(E)\varphi(B),\quad C(Q)=\varphi(Q^{t})C(E)\varphi(Q).$
@@
where $\mathbb{R}^{N}_{+}:=\left\{x=\left(x_{1},\cdots,x_{N}\right)\in\mathbb{R}^{N}\setminus\{0\}:x_{\alpha}\geq 0\ \mbox{for any}\ 1\leq\alpha\leq N\right\}$.
@@
With the given prerequisites, then $y$ is a maximum point of $f_{Q}\left(x\right)$ in the cone spanned by $\Delta_{\varepsilon}$ and an interior maximum point of $f_{Q}\left(x\right)$ in $\Delta_{\varepsilon}^{\gamma}$. Hence, applying the Lagrange Multiplier Method, there exist real numbers $a,b_{\gamma+1},\cdots,b_{N}$ such that
@@
(2.21) $\sum_{\beta=1}^{N}y_{\beta}\|[\hat{Q}_{\alpha},\hat{Q}_{\beta}]\|^{2}-\frac{1}{2}=\Big\{\begin{array}[]{ll}a&\text{if}\ \alpha\leq\gamma;\\ b_{\alpha}&\text{if}\ \alpha>\gamma.\end{array}$
@@
Hence $S\setminus(J\cup\{1\})\neq\varnothing$, and the last $\leq$” in (2.24) should be “<<” by the definition of $J$ and the positivity of $y_{\beta}$ for $\beta\in S\setminus(J\cup\{1\})$. $\Box$
@@
$y\in\Delta^{\gamma}:=\left\{x\in\Delta:x_{\alpha}>0\ \mbox{for}\ \alpha\leq\gamma,\ \mbox{and}\ x_{\beta}=0\ \mbox{for}\ \beta>\gamma\right\}$
@@
where $D=\operatorname{diag}(\lambda_{1},\ldots,\lambda_{q})$. In the proof of Lemma 2.3 with $J_{\alpha}=J$, the first line in (2.9) shows that $c_{\beta}=0$, $b_{\beta}$ is skew-symmetric and $b^{\beta}_{ij}=-b^{\beta}_{ji}=0$ for all $\beta\in J$ and $(i,j)\in\{(i,j):1\leq i<j\leq q\}\setminus I$; the third line in (2.9) shows that $b^{\beta}_{ij}=0$ for all $\beta\in S\setminus(J\cup\{1\})$ and $(i,j)\in I$. For the eigenvalues $\pm\lambda_{1},\ldots,\pm\lambda_{q}$ of $\hat{Q}_{1}$, because of the equality condition of Lemma 2.2, we have to consider the following cases:
@@ 3.3. Some isoparametric focal submanifolds satisfying Condition A
For an isoparametric hypersurface in a unit sphere with multiplicities $(m_{1},m_{2})$, there exist two focal submanifolds $M_{+}$ and $M_{-}$ of codimension $m_{1}+1$ and $m_{2}+1$, respectively. A point in a focal submanifold is said to satisfy Condition A if the kernels of all shape operators coincide [36]. The Condition A plays an important role in the research of isoparametric hypersurfaces in unit spheres with $4$ distinct principal curvatures. In [23], the subset $C_{A}$ of the points satisfying Condition A was determined for all the focal submanifolds. Thus, we know that the focal submanifolds satisfying Condition A everywhere are $M_{+}$ with multiplicities $(2,2)$, ($1,k)($$k\in\mathbb{N}^{+}$) and $M_{-}$ with multiplicities $(2,1)$, $(6,1)$, (4,3)(4,3)(definite). We are going to find austere submanifolds of type $\mathcal{Q}(p,q)$ in these focal submanifolds.
@@
$m$ 11 22 33 44 55 66 77 88 $\cdots~m+8$
$\delta(m)$ 11 22 44 44 88 88 88 88 $\cdots~16\delta(m)$
It follows that the focal submanifolds $M_{+}$ with multiplicities $(1,2k)$ are austere submanifolds of type $\mathcal{Q}(m_{2},m_{1}+m_{2})$. By interchanging the multiplicities $(m_{1},m_{2})$, we also obtain that the focal submanifolds $M_{-}$ with multiplicities $(2,1)$, $(6,1)$, (4,3)(4,3)(definite) are austere submanifolds of type $\mathcal{Q}(m_{1},m_{1}+m_{2})$. However, the focal submanifolds $M_{-}$ with multiplicities $(2,1)$, $(6,1)$ are congruent to focal submanifolds $M_{+}$ with multiplicities $(1,2)$, $(1,6)$.
@@
Let $M_{+}$ ((resp. $M_{-}$)) be the focal submanifold of an isoparametric hypersurface with $4$ distinct principal curvatures and codimension $m_{1}+1$ ((resp. $m_{2}+1$)) in a unit sphere.
@@
    The focal submanifold $M_{-}$ with multiplicities (4,3)(4,3)((definite)) is an austere submanifold of type $\mathcal{Q}(4,7)$.
@@ 4.2. Simons-type gap theorem
Since $\theta_{1},\cdots,\theta_{n}$ are orthonormal, we have $\theta_{j}=0$ for $j\geq 5$, and thus $\operatorname{dim}M=4$. Direct calculations show that for any $a,b\in\mathbb{R}$, the eigenvalues of $aC_{1}+bC_{2}$ are $\pm\sqrt{a^{2}+b^{2}}$ of multiplicity 2. In other words, $M$ has at most two distinct principal curvatures of multiplicity $2$ in every normal direction. Then it follows from [30, Theorem 1.5] that $M$ is given by the Veronese embeddings of complex projective plane. But the matrices corresponding to the second fundamental form of this Veronese embedding does not satisfy condition (4.4). (One can find an explicit description in [32, p. 560].) ∎